A child is running on a moving sidewalk in an airport. When she runs against the sidewalk's motion, she travels 60 in 10 seconds. When she runs with the sidewalk's motion, she travels the same distance in 5 seconds. What is the rate of the child on a still sidewalk and what is the rate of the moving sidewalk?
Answer by lwsshak3(11628)

A child is running on a moving sidewalk in an airport. When she runs against the sidewalk's motion, she travels 60 in 10 seconds. When she runs with the sidewalk's motion, she travels the same distance in 5 seconds. What is the rate of the child on a still sidewalk and what is the rate of the moving sidewalk?
**
let x=rate of child on a still sidewalk
let c=rate of moving sidewalk
x+c=rate of child walking with sidewalk's motion
x-c=rate of child walking against sidewalk's motion
distance/rate=travel time (same for both directions)
60/(x+c)=5
60/(x-c)=10
..
5x+5c=60
10x-10c=60
..
x+c=12
x-c=6
add
2x=18
x=9
c=12-x=3
rate of child on a still sidewalk=9 in/sec
rate of moving sidewalk=3 in/sec